Grade 11Algebra

If xlog y=log x then show that dy​/dx= y[1-logxlogy]/x(log2x)

Yamini
6 Years agoGrade 11

1 Answer

Aditya Gupta
6 Years ago
hey yamini
take log on both sides
log xlog y= log(log x)
logy*log = log(log x)
differentiate both sides wrt x
(1/y)*(y’)*log x + logy*1/x= 1/xlog x
so, y’=  dy​/dx= y[1 – logxlogy]/x(log2x)
kindly approve :)